OpenStudy (anonymous):

Simplify: (40a/b^2-5b+4) / (5a^2/2b-2)

OpenStudy (anonymous):

40a / (b-1)*(b-4) / 5a^2 / 2(b-1) 40 cancels with 5 leaving 8 a^2 cancels with a leaving a in denominator b-1 cancels with b-1 so left with 8/b-4 / a/2 invert and multiply 16/ab--4a

OpenStudy (anonymous):

can you understand ? @jchavez96

OpenStudy (anonymous):

Sorta , I'm working on it right now . @ashna

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

These are my options ashna , I'm having trouble figuring it out .

OpenStudy (anonymous):

@ashna

OpenStudy (anonymous):

4th option is right , you want to know how , or do you know ?

OpenStudy (anonymous):

@jchavez96

OpenStudy (anonymous):

Can you explain it to me please ?

OpenStudy (anonymous):

now take b^2-5b+4 and simplify it like u do spliting the middle term so....b^2-5b+4 b(b-4)-1(b-4)

OpenStudy (anonymous):

(b-1) (b-4)

OpenStudy (anonymous):

now {40a/(b-1)(b-4)}/{5a^2/2b-2}

OpenStudy (anonymous):

take the main denominater in reciprocal form then cancel 40a with 5a^2

OpenStudy (anonymous):

now 16b-16/(b-1)(b-4) = 16(b-1)/(b-1)(b-4)

OpenStudy (anonymous):

cancel b-1 u'll get opt4

OpenStudy (anonymous):

(i.e) = 16/b-4

OpenStudy (anonymous):

@jchavez96 understood now ?

OpenStudy (anonymous):

I see . . . Thank you !! What is the b=4 for ? The one with the equal sign crossed out ?

OpenStudy (anonymous):

@ashna

OpenStudy (anonymous):

dat's b not equal to 4 just substute b as 4 u wont get correct answer

OpenStudy (anonymous):

understood @jchavez96

OpenStudy (anonymous):

Yes ! Thank you @ashna ! I really appreciate it !

OpenStudy (anonymous):

no mention :)